meganmoney543:

Given: x + 2y = -6. Solve for x. What is the coefficient of y ?

AZ:

To solve for 'x', you have to get it all by itself. To do that, we have to move 2y to the other side. What do you get if you subtract 2y on both sides?

smackz:

1 attachment
AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @smackz \(\color{#0cbb34}{\text{End of Quote}}\) You solved for y not x. Also it would be -2y

smackz:

oop mb

AZ:

To answer the second part of your question: The coefficient is the number in front of the variable (or the letter) For example, 20a a is our variable 20 is the coefficient So what is the coefficient of 2y?

meganmoney543:

i jus need help

AZ:

x + 2y = -6 Subtract 2y on both sides x + 2y - 2y = -6 - 2y can you simplify it? What is 2y - 2y = ?

boredfr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @meganmoney543 i jus need help \(\color{#0cbb34}{\text{End of Quote}}\) He helped you.Restating what @AZ said, what is the coefficient of 2y?

meganmoney543:

nvm

meganmoney543:

my bf helped me

boredfr:

Then you honestly didn't want help, you just wanted the answer.